# Efficient computation of stationary measures and the Lyapunov Landscape for families random dynamical systems with smooth additive noise

### v1

@@ 7. Fourier theory and Galerkin approximation
We only give a short proof of item 2. Since $\{a_{k}\}_{k\in\mathbb{Z}}$ is in $\ell^{\exp}_{\sigma}$, we have
@@
This is the Fourier transform on $\mathbb{R}$ of a Gaussian [36, Theorem 1.4, Chapter 5]
@@
$D[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}dz=e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}.$
@@
$D_{\sigma}[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}\,dz=e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}.$
@@
    The image of $D_{\sigma}$ lies in the space of exponentially decaying sequences
    $\ell^{\exp}_{\sigma}:=\left\{(a_{k})_{k\in\mathbb{Z}}\;\middle|\;\exists C>0\text{ such that }|a_{k}|\leq Ce^{-\frac{\sigma^{2}|k|}{2}}\text{ for all }k\in\mathbb{Z}\right\}.$
@@ 8. Fourier Approximation
Theorem 8.4 (Tail Estimate for $N_{\sigma}$).
Let $\Gamma_{\sigma,k}:=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}$. Then:
$\|N_{\sigma}(1-\pi_{k})\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k},\quad\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k}.$
Sketch of Proof.
Let $f\in L^{1}([-1,1])$. By Parseval and the exponential decay of the diagonal entries $D[k,k]$ (see Lemma 7.7), we have
$\|N_{\sigma}(1-\pi_{k})f\|_{L^{2}}^{2}=\sum_{|j|>k}|D[j,j]|^{2}|\mathcal{F}(f)[j]|^{2}\leq\Gamma_{\sigma,k}^{2}\|f\|_{L^{1}}^{2}.$
@@
Follows directly from the identities $P_{\sigma}=N_{\sigma}P$ and $\pi_{k}^{2}=\pi_{k}$. ∎
@@
Lemma 8.6 (Norm Bound for Projected Operator).
The projected operator satisfies
$\|\pi_{k}N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma_{\sigma,k}.$
@@
We use the identity $\pi_{k}N_{\sigma}=N_{\sigma}\pi_{k}$, which follows from the commutation relations between $D_{\sigma}$, $\widetilde{\pi}_{k}$, and $\mathcal{F}$ (see Lemma 8.3). Then
$\|\pi_{k}N_{\sigma}f\|_{L^{1}}$ $\leq\|N_{\sigma}f\|_{L^{1}}+\|(1-\pi_{k})N_{\sigma}f\|_{L^{1}}$
$\leq\|f\|_{L^{1}}+\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{2}}\|f\|_{L^{1}}$
$\leq(1+\Gamma_{\sigma,k})\|f\|_{L^{1}}.$
@@
$\|\pi_{k}f_{\sigma}\|_{L^{1}}=\|\pi_{k}N_{\sigma}Pf_{\sigma}\|_{L^{1}}\leq(1+\Gamma_{\sigma,k})\|f_{\sigma}\|_{L^{1}},$
where $\Gamma_{\sigma,k}:=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}$.
@@
Note that if $f\in\mathcal{U}_{0}$, the space of zero-average functions in $L^{2}([-1,1])$, then
@@
Let $f_{\sigma}$ be the unique fixed point of $P_{\sigma}$, and let $P_{\sigma,k}$ be its Galerkin approximation. Suppose there exist $n\in\mathbb{N}$ and $C_{n}>0$ such that
$\|P_{\sigma,k}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C_{n}<1,$
and that for $0\leq i<n$, we have bounds $1\leq\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C_{i}$. Let $g$ be a trigonometric polynomial of degree at most $k$ such that $\|P_{\sigma,k}g-g\|_{L^{2}}\leq\epsilon$. Then
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left((1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}+\epsilon\right).$
@@
All constants involved in the theorem can be rigorously computed or estimated with the aid of a computer. In particular, the contraction bound $C_{n}$, the intermediate norms $C_{i}$, and the residual $\|P_{\sigma,k}g-g\|_{L^{2}}$ are all computable. The choice of the approximation $g$ is crucial: we typically use a numerically computed fixed point of $P_{\sigma,k}$, post-processed to enforce symmetry and improve approximation of the zero-frequency component.
@@
$\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|\leq\|P_{\sigma,k}^{n}\|^{q}\cdot\|P_{\sigma,k}^{r}\|\leq C_{n}^{q}C_{r}.$
@@
$\|P_{\sigma,k}^{i}\|\leq C_{n}^{\lfloor i/n\rfloor}\cdot\max_{r<n}C_{r}\leq C_{n}^{i/n-1}\cdot\max_{r<n}C_{r},$
@@
By Lemma 8.12(2), $\|P_{\sigma,k}^{n}v\|\to 0$ as $n\to\infty$. So:
@@
Let $\rho_{\sigma}(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}}$ be the Gaussian density, and let $\tau_{*}\rho_{\sigma}(x):=\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)$ be its $2$-periodic summation. Then
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}=\frac{1}{2\sigma\sqrt{\pi}}.$
@@
We prove both directions of the identity.
(1) Lower bound: By orthogonality of the supports (modulo period shifts),
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\int_{-1}^{1}\left|\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)\right|^{2}dx\geq\sum_{j\in\mathbb{Z}}\int_{-1}^{1}|\rho_{\sigma}(x+2j)|^{2}dx=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}.$
(2) Upper bound: Parseval’s identity for Fourier series on [1,1][-1,1] gives:
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=2\sum_{k\in\mathbb{Z}}|\mathcal{F}(\tau_{*}\rho_{\sigma})[k]|^{2}.$
The Fourier coefficients of $\tau_{*}\rho_{\sigma}$ are computed by extending the integral to $\mathbb{R}$:
$\mathcal{F}(\tau_{*}\rho_{\sigma})[k]=\frac{1}{2}\int_{\mathbb{R}}\rho_{\sigma}(x)e^{-k\pi ix}dx=\frac{1}{2}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
Hence,
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}.$
(3) Classical integral identity: It is well-known [34] that
$\int_{-\infty}^{+\infty}e^{-ax^{2}}dx=\sqrt{\frac{\pi}{a}}\quad\text{for }a>0.$
Applying this with $a=1/\sigma^{2}$,
$\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}=\int_{-\infty}^{\infty}\left(\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}}\right)^{2}dx=\frac{1}{\sigma^{2}(2\pi)}\int_{-\infty}^{\infty}e^{-\frac{x^{2}}{\sigma^{2}}}dx=\frac{1}{2\sigma\sqrt{\pi}}.$
This completes the proof. ∎
@@
$P_{\sigma,k}f_{\sigma}-f_{\sigma}=(\pi_{k}N_{\sigma}\pi_{k}P\pi_{k}-N_{\sigma}P)f_{\sigma}=(1-\pi_{k})N_{\sigma}P\pi_{k}f_{\sigma}+N_{\sigma}P(1-\pi_{k})f_{\sigma}.$
@@
$\|(1-\pi_{k})N_{\sigma}P\pi_{k}f_{\sigma}\|_{L^{2}}\leq\Gamma_{\sigma,k}\|\pi_{k}f_{\sigma}\|_{L^{1}}\leq\Gamma_{\sigma,k}(1+\Gamma_{\sigma,k})\|f_{\sigma}\|_{L^{1}}.$
For the second term, using Lemma 8.14 and Young’s inequality:
$\|N_{\sigma}P(1-\pi_{k})f_{\sigma}\|_{L^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}\cdot\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}.$
@@
$\delta=\|P_{\sigma,k}f_{\sigma}-f_{\sigma}\|_{L^{2}}\leq(1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}.$

### v2

@@ 7. Fourier theory and Galerkin approximation
Proofs of items1 and 3 can be found in Theorem 9.6 and Theorem 9.13 of [43], respectively. We only give a short proof of item 2. Since $\{a_{k}\}_{k\in\mathbb{Z}}$ is in $\ell^{\exp}_{\sigma}$, we have
@@
This is the Fourier transform on $\mathbb{R}$ of a Gaussian [45, Theorem 1.4, Chapter 5]
@@
$D[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}dz=e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
@@
$D_{\sigma}[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}\,dz=e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
@@
    The image of $D_{\sigma}$ lies in the space of exponentially decaying sequences $\ell^{\exp}_{\sigma}$.
@@ 8. Fourier Approximation
Theorem 8.4.
Let $\pi_{k}$ denote the Fourier projection onto modes $|j|\leq k$ on [1,1][-1,1], and let $N_{\sigma}$ be the Gaussian noise operator. Set
$\Gamma_{\sigma,k}:=\sqrt{\frac{1}{\sigma\sqrt{\pi}}\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)}\;e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
Then
$\|N_{\sigma}(1-\pi_{k})\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k},\qquad\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k}.$
Proof.
From Parseval’s identity, we have
(8) $\|g\|_{L^{2}([-1,1])}^{2}=2\sum_{j\in\mathbb{Z}}|\mathcal{F}(g)[j]|^{2}.$
Moreover, $N_{\sigma}$ is a Fourier multiplier:
$\mathcal{F}(N_{\sigma}f)[j]=D[j,j]\,\mathcal{F}(f)[j],$
and $\pi_{k}$ is also a Fourier multiplier. Hence $N_{\sigma}$ and $\pi_{k}$ commute, and $N_{\sigma}(1-\pi_{k})=(1-\pi_{k})N_{\sigma}$, so it suffices to bound one of them.
Let $f\in L^{1}([-1,1])$. Since $(1-\pi_{k})$ removes modes $|j|\leq k$,
$\|N_{\sigma}(1-\pi_{k})f\|_{L^{2}}^{2}=2\sum_{|j|>k}|D[j,j]|^{2}\,|\mathcal{F}(f)[j]|^{2}.$
Using $|\mathcal{F}(f)[j]|\leq\frac{1}{2}\|f\|_{L^{1}}$ (Lemma 7.5(1)), we obtain
$\|N_{\sigma}(1-\pi_{k})f\|_{L^{2}}^{2}\leq 2\sum_{|j|>k}e^{-\sigma^{2}\pi^{2}j^{2}}\,\frac{1}{4}\|f\|_{L^{1}}^{2}=\frac{1}{2}\Bigl(\sum_{|j|>k}e^{-\sigma^{2}\pi^{2}j^{2}}\Bigr)\|f\|_{L^{1}}^{2}.$
To bound the tail sum, write $j=k+m$ with $m\geq 1$ and note $(k+m)^{2}\geq k^{2}+m^{2}$, hence
$\sum_{j\geq k+1}e^{-\sigma^{2}\pi^{2}j^{2}}\leq e^{-\sigma^{2}\pi^{2}k^{2}}\sum_{m\geq 1}e^{-\sigma^{2}\pi^{2}m^{2}}\leq e^{-\sigma^{2}\pi^{2}k^{2}}\sum_{m\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}m^{2}}.$
Therefore
$\sum_{|j|>k}e^{-\sigma^{2}\pi^{2}j^{2}}\leq 2e^{-\sigma^{2}\pi^{2}k^{2}}\sum_{m\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}m^{2}}.$
By Jacobi/Poisson summation,
$\sum_{m\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}m^{2}}=\frac{1}{\sigma\sqrt{\pi}}\sum_{n\in\mathbb{Z}}e^{-n^{2}/\sigma^{2}}\leq\frac{1}{\sigma\sqrt{\pi}}\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr),$
using the bound from Remark 8.16. Combining the previous estimates gives
$\|N_{\sigma}(1-\pi_{k})f\|_{L^{2}}^{2}\leq\frac{1}{2}\cdot 2e^{-\sigma^{2}\pi^{2}k^{2}}\cdot\frac{1}{\sigma\sqrt{\pi}}\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)\,\|f\|_{L^{1}}^{2},$
and taking square roots yields the stated bound. ∎
Remark 8.5.
The factor $\coth\!\bigl(\frac{1}{2\sigma^{2}}\bigr)$ controls the wrap–around effect introduced by periodizing the Gaussian kernel (cf. Lemma 8.17 and Remark 8.16).
@@
Follows directly from the identities $P_{\sigma}=N_{\sigma}P$, $\pi_{k}^{2}=\pi_{k}$ and $\pi_{k}N_{\sigma}=N_{\sigma}\pi_{k}$; the latter follows from Proposition 7.6 and the commutation relations between $D_{\sigma}$, $\widetilde{\pi}_{k}$, and $\mathcal{F}$ (see Lemma 8.3). ∎
@@
Lemma 8.7.
Let $N_{\sigma}$ be the Gaussian noise operator with periodic boundary condition, and let $\pi_{k}$ be the Fourier projection onto modes $|j|\leq k$. Then for every $f\in L^{1}([-1,1])$,
(9) $\|(1-\pi_{k})N_{\sigma}f\|_{L^{\infty}([-1,1])}$ $\leq\frac{1}{2}\|f\|_{L^{1}([-1,1])}\sum_{|j|>k}e^{-\frac{\sigma^{2}\pi^{2}j^{2}}{2}},$
(10) $\|(1-\pi_{k})N_{\sigma}f\|_{L^{1}([-1,1])}$ $\leq\|f\|_{L^{1}([-1,1])}\sum_{|j|>k}e^{-\frac{\sigma^{2}\pi^{2}j^{2}}{2}}.$
In particular,
$\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{1}}\leq\Gamma^{(1)}_{\sigma,k},\qquad\Gamma^{(1)}_{\sigma,k}:=\sum_{|j|>k}e^{-\frac{\sigma^{2}\pi^{2}j^{2}}{2}}.$
Moreover, letting $a:=\sigma^{2}\pi^{2}/2$, the tail admits the explicit bound
(11) $\Gamma^{(1)}_{\sigma,k}\leq\frac{2}{\sigma^{2}\pi^{2}\,k}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}\qquad(k\geq 1).$
@@
Since $N_{\sigma}$ is a Fourier multiplier,
$\mathcal{F}(N_{\sigma}f)[j]=D[j,j]\mathcal{F}(f)[j],$
Therefore
$(1-\pi_{k})N_{\sigma}f(x)=\sum_{|j|>k}D[j,j]\mathcal{F}(f)[j]\,e^{j\pi ix}.$
Using $|\mathcal{F}(f)[j]|\leq\frac{1}{2}\|f\|_{L^{1}}$, we obtain for each $x\in[-1,1]$,
$|(1-\pi_{k})N_{\sigma}f(x)|\leq\sum_{|j|>k}D[j,j]|\mathcal{F}(f)[j]|\leq\frac{1}{2}\|f\|_{L^{1}}\sum_{|j|>k}e^{-\frac{\sigma^{2}\pi^{2}j^{2}}{2}},$
which yields (9). Since m([1,1])=2m([-1,1])=2, we also have $\|g\|_{L^{1}([-1,1])}\leq 2\|g\|_{L^{\infty}([-1,1])}$, hence (10).
For (11), write $\Gamma^{(1)}_{\sigma,k}=\displaystyle 2\sum_{j\geq k+1}e^{-aj^{2}}$ and bound the tail by an integral:
$\sum_{j\geq k+1}e^{-aj^{2}}\leq\int_{k}^{\infty}e^{-ax^{2}}\,dx\leq\frac{e^{-ak^{2}}}{2ak},$
which gives $\Gamma^{(1)}_{\sigma,k}\leq\frac{1}{ak}e^{-ak^{2}}=\frac{2}{\sigma^{2}\pi^{2}k}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}$. ∎
Corollary 8.8.
With $\Gamma^{(1)}_{\sigma,k}$ as in Lemma 8.7, one has
$\|\pi_{k}N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma^{(1)}_{\sigma,k}$.
Proof.
Since $\pi_{k}=\textrm{Id}-(\textrm{Id}-\pi_{k})$ and $\|N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1$ (Lemma 6.2),
$\|\pi_{k}N_{\sigma}\|_{L^{1}\to L^{1}}\leq\|N_{\sigma}\|_{L^{1}\to L^{1}}+\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma^{(1)}_{\sigma,k}$.
@@
$\|\pi_{k}f_{\sigma}\|_{L^{1}}=\|\pi_{k}N_{\sigma}Pf_{\sigma}\|_{L^{1}}\leq(1+\Gamma^{(1)}_{\sigma,k})\|f_{\sigma}\|_{L^{1}}.$
@@
Note that if $f\in\mathcal{U}_{0}$, the space of average 0 functions in $L^{2}([-1,1])$, then
@@
Let $f_{\sigma}$ be the unique fixed point of $P_{\sigma}$, and let $P_{\sigma,k}$ be its Galerkin approximation. For $i\in\mathbb{N}_{0}$, set $C_{i}:=\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}$, and suppose that $C_{n}<1$ for some $n\in\mathbb{N}$. If $g$ be a trigonometric polynomial of degree at most $k$ such that $\|P_{\sigma,k}g-g\|_{L^{2}}\leq\epsilon$. Then
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left((1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma^{(1)}_{\sigma,k}\|f_{\sigma}\|_{L^{1}}+\epsilon\right).$
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left(\delta+\epsilon\right).$
where
$\delta\leq\left(\Gamma_{\sigma,k}(1+\Gamma_{\sigma,k}^{(1)})+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}\sqrt{\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)}\cdot\Gamma^{(1)}_{\sigma,k}\right)\|f_{\sigma}\|_{L^{1}}$
@@
All constants involved in the theorem can be rigorously computed or estimated with the aid of a computer. In particular, the contraction bound $C_{n}$, the intermediate norms $C_{i}$, and the residual $\|P_{\sigma,k}g-g\|_{L^{2}}$ are all computable. The choice of the approximation $g$ is crucial: we typically use a numerically computed fixed point of $P_{\sigma,k}$, and enforce symmetry in its truncated Fourier representation to ensure that the resulting trigonometric polynomial is real. An approximation $g$ choice in this way will be referred to as a symmetrized trigonometric polynomial approximating $f_{\sigma}$
@@
$\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq\|P_{\sigma,k}^{n}\|_{L^{2}\rightarrow L^{2}}^{q}\cdot\|P_{\sigma,k}^{r}\|_{L^{2}\rightarrow L^{2}}\leq C_{n}^{q}C_{r}.$
@@
$\|P_{\sigma,k}^{i}\|_{L^{2}\rightarrow L^{2}}\leq C_{n}^{\lfloor i/n\rfloor}\cdot\max_{r<n}C_{r}\leq C_{n}^{i/n-1}\cdot\max_{r<n}C_{r},$
@@
By Lemma 8.14(2), $\|P_{\sigma,k}^{n}v\|_{L^{2}}\to 0$ as $n\to\infty$. So:
@@
Remark 8.16.
For $\sigma>0$ we have the elementary estimate $n^{2}\geq n$ for all $n\geq 1$, hence
$\sum_{n\geq 1}e^{-n^{2}/\sigma^{2}}\leq\sum_{n\geq 1}e^{-n/\sigma^{2}}=\frac{e^{-1/\sigma^{2}}}{1-e^{-1/\sigma^{2}}}.$
Therefore
$1+2\sum_{n\geq 1}e^{-n^{2}/\sigma^{2}}\leq 1+\frac{2e^{-1/\sigma^{2}}}{1-e^{-1/\sigma^{2}}}=\frac{1+e^{-1/\sigma^{2}}}{1-e^{-1/\sigma^{2}}}=\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr).$
@@
Let $\rho_{\sigma}(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})}$ and let
$(\tau_{*}\rho_{\sigma})(x):=\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j),\qquad x\in[-1,1],$
be its 22–periodization. Then $\tau_{*}\rho_{\sigma}\in L^{2}([-1,1])$ and
(12) $\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\frac{1}{2}\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}}.$
Moreover,
(13) $\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}=\frac{1}{2\sigma\sqrt{\pi}},$
and, since $\rho_{\sigma}>0$ everywhere,
(14) $\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}>\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}.$
Equivalently, using the Poisson summation identity,
(15) $\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\frac{1}{2\sigma\sqrt{\pi}}\sum_{n\in\mathbb{Z}}e^{-n^{2}/\sigma^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr).$
@@
For $f=\tau_{*}\rho_{\sigma}$, we compute for each $k\in\mathbb{Z}$:
$\mathcal{F}(\tau_{*}\rho_{\sigma})[k]$ $=\frac{1}{2}\int_{-1}^{1}\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)\,e^{-k\pi ix}\,dx$
$=\frac{1}{2}\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\rho_{\sigma}(x+2j)\,e^{-k\pi ix}\,dx=\frac{1}{2}\int_{\mathbb{R}}\rho_{\sigma}(u)\,e^{-k\pi iu}\,du,$
where we used Tonelli/Fubini (absolute convergence) and the change of variables $u=x+2j$. Since the Fourier transform of a Gaussian is Gaussian,
$\int_{\mathbb{R}}\rho_{\sigma}(u)\,e^{-k\pi iu}\,du=e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}},$
hence
$\mathcal{F}(\tau_{*}\rho_{\sigma})[k]=\frac{1}{2}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
Plugging this into (8) yields
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=2\sum_{k\in\mathbb{Z}}\Bigl|\frac{1}{2}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}\Bigr|^{2}=\frac{1}{2}\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}},$
which is (12).
For (13), we compute
$\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}=\int_{\mathbb{R}}\Bigl(\frac{1}{\sigma\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})}\Bigr)^{2}\,dx=\frac{1}{2\pi\sigma^{2}}\int_{\mathbb{R}}e^{-x^{2}/\sigma^{2}}\,dx=\frac{1}{2\sigma\sqrt{\pi}}.$
Finally, (14) follows immediately from (12) and Jabobi/Poisson summation:
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\frac{1}{2}\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}}=\frac{1}{2}\left(\frac{1}{\sigma\sqrt{\pi}}\displaystyle\sum_{n\in\mathbb{Z}}e^{-n^{2}/\sigma^{2}}\right)>\frac{1}{2\sigma\sqrt{\pi}}=\|\rho_{\sigma}\|^{2}_{L^{2}(\mathbb{R})}$
i.e. the extra $n\neq 0$ terms are strictly positive. The alternative expression (15) follows from the Poisson summation formula applied to $\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}}$. ∎
@@
$P_{\sigma,k}f_{\sigma}-f_{\sigma}=(\pi_{k}N_{\sigma}\pi_{k}P\pi_{k}-N_{\sigma}P)f_{\sigma}=(\pi_{k}-1)N_{\sigma}P\pi_{k}f_{\sigma}+N_{\sigma}P(\pi_{k}-1)f_{\sigma}.$
@@
$\|(1-\pi_{k})N_{\sigma}P\pi_{k}f_{\sigma}\|_{L^{2}}\leq\Gamma_{\sigma,k}\|\pi_{k}f_{\sigma}\|_{L^{1}}\leq\Gamma_{\sigma,k}(1+\Gamma^{(1)}_{\sigma,k})\|f_{\sigma}\|_{L^{1}}.$
For the second term, using Lemma 8.17, Remark 8.16 Lemma 8.7 and Young’s inequality:
$\|N_{\sigma}P(1-\pi_{k})f_{\sigma}\|_{L^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}\sqrt{\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)}\cdot\Gamma^{(1)}_{\sigma,k}\|f_{\sigma}\|_{L^{1}}.$
@@
$\delta=\|P_{\sigma,k}f_{\sigma}-f_{\sigma}\|_{L^{2}}\leq\left(\Gamma_{\sigma,k}(1+\Gamma_{\sigma,k}^{(1)})+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}\sqrt{\coth\!\Bigl(\frac{1}{2\sigma^{2}}\Bigr)}\cdot\Gamma^{(1)}_{\sigma,k}\right)\|f_{\sigma}\|_{L^{1}}$